Liquid N2O3 is dark blue at low temperatures, but the color fades and becomes greenish at higher temperatures as the compound decomposes to NO and NO2. At 25 degree C, a value of Kp = 1.91 has been established for this decomposition. If 0236 moles of N2O3 are placed in a 1.52-L vessel at 25 degree C, calculate the equilibrium partial pressures of N2O3(g), NO2(g) and NO(g).
1
Expert's answer
2017-01-30T14:18:13-0500
N2O3 = NO2 + NO

First, the mol of each component can be calculated from the equation [NO2][NO]/[N2O3] = Kp. Thus, C(N2O3) = 0.026 mol, C(NO2) = 0.210 mol and C(NO) = 0.210 mol.
According to a perfect gas state equation p = n*R*T/V. Thus, p(N2O3) = 42.4 kPa, p(NO2) = 342 kPa and p(NO) = 342 kPa.